OpenStudy (anonymous):

solve the equation 5 over 4 minus 1 over x minus 2 equals -4 over 4x minus. use complete sentences to explain how you solved the equation

OpenStudy (anonymous):

Could you please make this more clear as to which numbers are in the numerators and denominators of each equations? And 4x minus what?

OpenStudy (anonymous):

5/4 - 1/x-2 = -4/4x-8 sorry I typed it in wrong

OpenStudy (anonymous):

The easiest thing to do here is first factor the right side of the equation. 5/4-1/(x-2)=-4/(4(x-2)) We can then multiply both sides by x-2 5(x-2)/4-1=-4/4 5(x-2)/4-1=-1 Now solve like any other linear equation. 5(x-2)/4=0 5(x-2)=0 x-2=0 x=2 But, because 2 makes the denominator of both equations zero, we can say the system has no real solutions.

OpenStudy (anonymous):

Thanks :)